OpenStudy (anonymous):

Find the point on the terminal side of θ = -3pie/4 that has an x coordinate of -1. Show full work.

OpenStudy (yttrium):

Maybe it's time for you to do this work. :)) Just show me what you can. I'll support you.

OpenStudy (debbieg):

Remember that for ANY point on the terminal side of \(-3\pi/4\), \[\large \tan \left(-\dfrac{3\pi}{4}\right)=\frac{ y }{ x }\] So if you know that tangent value (you should, it's a standard unit circle angle), then just set that = to y/(-1) and find the y coordinate.

OpenStudy (anonymous):

Should my answer me a fraction or a decimal?

OpenStudy (anonymous):

Be*

OpenStudy (anonymous):

is it y = 13.42?

OpenStudy (anonymous):

@DebbieG ?

OpenStudy (anonymous):

@PFEH.1999 can you maybe help me with this?

OpenStudy (anonymous):

Wow really? Thanks everyone!

OpenStudy (debbieg):

No, that's not the y value. Can you show how you got that?

OpenStudy (anonymous):

Its not? I set it = to y/ -1 and cross multiplied

OpenStudy (debbieg):

you set WHAT = y/(-1)? \(\large \tan \left(-\dfrac{3\pi}{4}\right)=\frac{ y }{ x }\) Should be the TANGENT VALUE that you are setting = y/(-1). What is \(\large \tan \left(-\dfrac{3\pi}{4}\right)\)?

OpenStudy (anonymous):

Im not sure how to find the Tangent value of that

OpenStudy (debbieg):

ok, well, that's a critical part here. You are doing function values with unit circle angles, so how have you been taught to do them? Have you learned the two "special" right triangles and the side ratios? 

OpenStudy (debbieg):

Or, are you supposed to refer to a unit circle?

OpenStudy (debbieg):

Since you are given this problem, it's clear that you are - somehow - supposed to know (or be able to find) the tangent value.

OpenStudy (debbieg):

Another approach to finding y, if you understand the triangle approach would be this: